# AReferences

### v1

@@ 5 Related Work
Influence estimation has been a critical area of research in understanding the impact of individual training samples on machine learning models [36, 31, 22, 43, 6]. Early work by Koh and Liang [16], Agarwal et al. [1] proposed second-order Hessian-based methods to approximate the effect of a training sample. However, approximating a Hessian inversion becomes computationally prohibitive for large-scale datasets and modern models containing billions of parameters. To address this issue, some studies proposed first order approaches for influence estimation [32, 31]. However, even with first-order methods, scaling to large datasets still encounter storage challenges. For example, storing the gradient of a 2B diffusion model for 10,000 data samples across 10 timesteps requires over 700 TB of storage.
To reduce the storage and computational demands, some studies leverage dimension reduction techniques [33, 31, 30, 11, 14], such as random projection. However, while random projection can substantially reduce the dimension of gradient vector, the projection matrix itself becomes a scalability bottleneck in large models. For instance, in a model with 2B parameters, a projection matrix mapping gradients to a compressed dimension of 32,768 would require over 500 GB of storage. These constraints highlight the need for more efficient and scalable approaches.
@@ 6 Conclusion
In this paper, we introduce DMin, a scalable framework for estimating the influence of training data samples on images generated by diffusion models. The proposed DMin substantially reduces storage requirements from 339.39 TB to only 726 MB for SD 3 with full parameters, and retrieves the top-$k$ most influential training sample in 1 second by KNN. This confirms that the proposed DMin can be scaled to the diffusion models with billions of parameters. Our empirical results confirm DMin’s effectiveness and efficiency.
@@ 7 Experimental Settings
Implementation Details e provide an open-source PyTorch implementation with multiprocessing support [^2] https://anonymous.4open.science/r/DMin. We leverage Huggingface, Accelerate, Transformers, Diffusers and Peft in our implementation.
Experimental Environments Our experiments are conducted on four different types of servers: (1) Servers running Red Hat Enterprise Linux 7.8, equipped with Intel(R) Xeon(R) Platinum 8358 processors (2.60GHz) with 32 cores, 64 threads, 4 A100 80G GPUs, and 1TB of memory. (3) Servers running Red Hat Enterprise Linux 7.8, containing Intel(R) Xeon(R) Gold 6226R CPUs @ 2.90GHz with 16 cores, 32 threads, 2 A100 40G GPUs, and 754GB of memory. (4) A server running Ubuntu 20.04.6 LTS, featuring 2 H100 GPUs, dual Intel(R) Xeon(R) Gold 6438N processors (3.60GHz) with 32 cores, 64 threads, and 1.48TB of memory. To ensure a fair comparison, all experiments measuring time cost and memory consumption are conducted on server 1, while other experiments are distributed across the different server types.
@@ 7.2 Datasets
200 Flowers 4 0.9056 0.9046 0.9051 0.9062 0.9062 0.9062
8 0.9062 0.9062 0.9062 0.9062 0.9062 0.9062
16 0.9062 0.9062 0.9062 0.9062 0.9062 0.9062
32 0.9062 0.9062 0.9062 0.9062 0.9062 0.9062
48 0.9062 0.9062 0.9062 0.9062 0.9062 0.9062
Lego Sets 4 0.4835 0.5647 0.5647 0.5647 0.5647 0.5647
8 0.5624 0.5624 0.5624 0.5624 0.5624 0.5624
16 0.5624 0.5624 0.5624 0.5624 0.5624 0.5624
32 0.5624 0.5624 0.5624 0.5624 0.5624 0.5624
48 0.5624 0.5624 0.5624 0.5624 0.5624 0.5624
Magic Cards 4 0.9826 0.9842 0.9827 0.9827 0.9827 0.9827
8 0.9825 0.9825 0.9824 0.9824 0.9824 0.9824
16 0.9824 0.9824 0.9824 0.9824 0.9824 0.9824
32 0.9824 0.9824 0.9824 0.9824 0.9824 0.9824
48 0.9825 0.9824 0.9824 0.9824 0.9824 0.9824
1000 Flowers 4 0.9062 0.9056 0.9062 0.9062 0.9062 0.9062
8 0.9062 0.9062 0.9062 0.9062 0.9062 0.9062
16 0.9062 0.9062 0.9062 0.9062 0.9062 0.9062
32 0.9062 0.9062 0.9062 0.9062 0.9062 0.9062
48 0.9062 0.9062 0.9062 0.9062 0.9062 0.9062
Lego Sets 4 0.52 0.5624 0.5635 0.5635 0.5635 0.5635
8 0.5624 0.5624 0.5624 0.5624 0.5624 0.5624
16 0.5624 0.5624 0.5624 0.5624 0.5624 0.5624
32 0.5624 0.5624 0.5624 0.5624 0.5624 0.5624
48 0.5624 0.5624 0.5624 0.5624 0.5624 0.5624
Magic Cards 4 0.9825 0.9824 0.9824 0.9824 0.9824 0.9824
8 0.9824 0.9824 0.9824 0.9824 0.9824 0.9824
16 0.9824 0.9824 0.9824 0.9824 0.9824 0.9824
32 0.9824 0.9824 0.9824 0.9824 0.9824 0.9824
48 0.9824 0.9824 0.9824 0.9824 0.9824 0.9824
@@
SD 1.4 (LoRA) 0.001 88 150 $512\times 512$ 4 8 [to_k, to_q, to_v, to_out.0] float32
SD 3 Medium (LoRA) 0.001 8 150 $512\times 512$ 4 8 [to_k, to_q, to_v, to_out.0] float32
SD 3 Medium (Full) 0.0001 8 150 $512\times 512$ float32
@@
Model Method Flowers Lego Sets Magic Cards
SD 1.4 (LoRA) Ours ($v=2^{12}$, KNN, HNSW) 0.9846 0.9538 0.7513 0.5290 0.6941 0.6176 0.2400 0.1412 0.9604 0.9463 0.9356 0.9364
Ours ($v=2^{16}$, KNN, HNSW) 0.9846 0.9564 0.7569 0.5492 0.7294 0.6235 0.2506 0.1441 0.987 0.984 0.9731 0.9724
Ours ($v=2^{12}$, KNN, Faiss) 0.8000 0.7949 0.7031 0.3982 0.5294 0.5059 0.2612 0.1306 0.9133 0.9133 0.9114 0.9125
Ours ($v=2^{16}$, KNN, Faiss) 0.8667 0.8359 0.7395 0.4105 0.5176 0.4941 0.2494 0.1253 0.9837 0.9837 0.9826 0.9841
SD 3 Medium (LoRA) Ours ($v=2^{12}$, KNN, HNSW) 0.9487 0.9385 0.9062 0.8544 0.9294 0.8882 0.5624 0.3141 0.9919 0.9913 0.9824 0.9774
Ours ($v=2^{16}$, KNN, HNSW) 0.9487 0.9436 0.9123 0.8608 0.9412 0.9000 0.5553 0.3129 0.9919 0.9916 0.9844 0.9787
Ours ($v=2^{12}$, KNN, Faiss) 0.9231 0.9231 0.8595 0.5533 0.8471 0.8353 0.6776 0.3394 0.9919 0.9916 0.9885 0.9902
Ours ($v=2^{16}$, KNN, Faiss) 0.9487 0.9487 0.7887 0.3944 0.8824 0.8529 0.6812 0.3412 0.9919 0.9913 0.9837 0.9878
SD 3 Medium (Full) Ours ($v=2^{12}$, KNN, HNSW) 0.8923 0.8615 0.6805 0.4713 0.85 0.8125 0.4000 0.2275 0.9745 0.9663 0.9498 0.9413
Ours ($v=2^{16}$, KNN, HNSW) 0.8923 0.8615 0.6949 0.4836 0.8375 0.8187 0.3775 0.2181 0.9810 0.9671 0.9520 0.9435
Ours ($v=2^{12}$, KNN, Faiss) 0.8000 0.7744 0.5903 0.4297 0.8353 0.8294 0.5953 0.2976 0.9458 0.9298 0.8887 0.8986
Ours ($v=2^{16}$, KNN, Faiss) 0.8462 0.8385 0.6236 0.4564 0.8824 0.8824 0.5659 0.2829 0.9442 0.9339 0.8989 0.8971
Table 8: Comparison of average detection rates using the Faiss KNN algorithm across various models and subsets.
@@ 8 Ablation Study
To better understand the impact of key parameters on the performance of the HNSW implementation, we conducted an ablation study by varying the graph-related parameters $M$ and ef, as well as the construction parameter $\text{ef}_{\text{construction}}$. Table 6 summarizes the average detection rates across three subsets: Flowers, Lego Sets, and Magic Cards, under a range of settings.
@@
Across the three datasets, the Magic Cards consistently exhibited high detection rates, exceeding 98.2% in all configurations, indicating that it is less sensitive to parameter tuning. In contrast, the Lego Sets showed significant variability. For $\text{ef}=200$, the detection rate improved notably with higher values of $M$ (e.g., from 48.3% at $M=4$ to 56.2% at $M=8$), but beyond $\text{ef}_{\text{construction}}=200$, further increases in $\text{ef}_{\text{construction}}$ provided diminishing returns. This suggests that while denser graphs and more exhaustive index construction improve accuracy for complex datasets, the benefits plateau at a certain point. For the Flowers, the detection rates remained stable at approximately 90.6% across all parameter settings, indicating that this dataset is robust to variations in $M$ and $ef$.
@@ 8.1 Baselines
9 KNN Algorithms
In this experiments, we replaced HNSW with Faiss algorithm to evaluate its performance in detecting the top-$k$ most influential training samples. Faiss is a KNN algorithm leveraging optimized vector operations and inner product similarity. We report the comparison of average detection rates in Table 8, HNSW demonstrates competitive or superior detection rates compared to Faiss across datasets.
@@ 10 Supplemental Visualization for Conditional Diffusion Models
Figure 6: Examples of the top-25 most influential training data samples for the generated image (the 1-st row) on SD 3 Medium with LoRA, shown from high to low influence from left to right.

### v3

@@ 5 Related Work
Influence estimation has been a critical area of research in understanding the impact of individual training samples on machine learning models [32, 28, 37, 6]. Early work by Koh and Liang [16], Agarwal et al. [1] proposed second-order Hessian-based methods to approximate the effect of a training sample. However, approximating a Hessian inversion becomes computationally prohibitive for large-scale datasets and modern models containing billions of parameters. To address this issue, some studies proposed first order approaches for influence estimation [29, 28]. However, even with first-order methods, scaling to large datasets still encounter storage challenges. For example, storing the gradient of a 2B diffusion model for 10,000 data samples across 10 timesteps requires over 700 TB of storage.
To reduce the storage and computational demands, some studies leverage dimension reduction techniques [28, 27, 11, 14], such as random projection. However, while random projection can substantially reduce the dimension of gradient vector, the projection matrix itself becomes a scalability bottleneck in large models. For instance, in a model with 2B parameters, a projection matrix mapping gradients to a compressed dimension of 32,768 would require over 500 GB of storage. These constraints highlight the need for more efficient and scalable approaches.
@@ 6 Conclusion
In this paper, we introduce DMin, a scalable framework for estimating the influence of training data samples on images generated by diffusion models. The proposed DMin scales effectively to diffusion models with billions of paramters by substantially reducing storage requirements from hundreds of TBs to MBs or KBs for SD 3 Medium with full parameters. Additionally, DMin can retrieve the top-$k$ most influential training sample in 1 second by KNN, demonstrating the scalability of the proposed DMin. Our empirical results further confirm DMin’s effectiveness and efficiency.
@@
Implementation Details. We provide an open-source PyTorch implementation with multiprocessing support. We leverage Huggingface, Accelerate, Transformers, Diffusers and Peft in our implementation.
Experimental Environments. Our experiments are conducted on four different types of servers: (1) Servers running Red Hat Enterprise Linux 7.8, equipped with Intel(R) Xeon(R) Platinum 8358 processors (2.60GHz) with 32 cores, 64 threads, 4 A100 80G GPUs, and 1TB of memory. (3) Servers running Red Hat Enterprise Linux 7.8, containing Intel(R) Xeon(R) Gold 6226R CPUs @ 2.90GHz with 16 cores, 32 threads, 2 A100 40G GPUs, and 754GB of memory. (4) A server running Ubuntu 20.04.6 LTS, featuring 2 H100 GPUs, dual Intel(R) Xeon(R) Gold 6438N processors (3.60GHz) with 32 cores, 64 threads, and 1.48TB of memory. To ensure a fair comparison, all experiments measuring time cost and memory consumption are conducted on server 1, while other experiments are distributed across the different server types.
@@ 7.2 Datasets
200 Flowers 4 0.8405 0.8349 0.8405 0.8405 0.8405 0.8405
8 0.8410 0.8410 0.8415 0.8415 0.8415 0.8415
16 0.8410 0.8415 0.8415 0.8415 0.8415 0.8415
32 0.8410 0.8415 0.8415 0.8415 0.8415 0.8415
48 0.8410 0.8415 0.8415 0.8415 0.8415 0.8415
Lego Sets 4 0.2800 0.3035 0.3094 0.3094 0.3082 0.3082
8 0.3082 0.3082 0.3094 0.3094 0.3094 0.3094
16 0.3082 0.3082 0.3094 0.3082 0.3094 0.3094
32 0.3082 0.3082 0.3094 0.3094 0.3094 0.3094
48 0.3082 0.3082 0.3094 0.3094 0.3094 0.3094
Magic Cards 4 0.9770 0.9772 0.9772 0.9772 0.9772 0.9772
8 0.9772 0.9772 0.9772 0.9771 0.9771 0.9771
16 0.9771 0.9771 0.9771 0.9771 0.9771 0.9771
32 0.9771 0.9771 0.9771 0.9771 0.9771 0.9771
48 0.9771 0.9771 0.9771 0.9771 0.9771 0.9771
1000 Flowers 4 0.8415 0.8415 0.8415 0.8415 0.8415 0.8415
8 0.8415 0.8415 0.8415 0.8415 0.8415 0.8415
16 0.8415 0.8415 0.8415 0.8415 0.8415 0.8415
32 0.8415 0.8415 0.8415 0.8415 0.8415 0.8415
48 0.8415 0.8415 0.8415 0.8415 0.8415 0.8415
Lego Sets 4 0.2847 0.3035 0.3094 0.3094 0.3094 0.3094
8 0.3094 0.3094 0.3094 0.3094 0.3094 0.3094
16 0.3094 0.3094 0.3094 0.3094 0.3094 0.3094
32 0.3094 0.3094 0.3094 0.3094 0.3094 0.3094
48 0.3094 0.3094 0.3094 0.3094 0.3094 0.3094
Magic Cards 4 0.9770 0.9771 0.9771 0.9771 0.9771 0.9771
8 0.9771 0.9771 0.9771 0.9771 0.9771 0.9771
16 0.9771 0.9771 0.9771 0.9771 0.9771 0.9771
32 0.9771 0.9771 0.9771 0.9771 0.9771 0.9771
48 0.9771 0.9771 0.9771 0.9771 0.9771 0.9771
@@
SD 1.4 (LoRA) 0.001 64 150 $512\times 512$ 4 8 [to_k, to_q, to_v, to_out.0] float32
SD 3 Medium (LoRA) 0.001 64 150 $512\times 512$ 4 8 [to_k, to_q, to_v, to_out.0] float32
SD 3 Medium (Full) 0.0001 64 150 $512\times 512$ float32
@@ 8 Ablation Study
To better understand the impact of key parameters on the performance of the HNSW implementation, we conducted an ablation study by varying the graph-related parameters $M$ and ef, as well as the construction parameter $\text{ef}_{\text{construction}}$. Table 6 summarizes the average detection rates across three subsets: Flowers, Lego Sets, and Magic Cards, under a range of settings on SD 3 Medium with LoRA ($v=2^{12}$).
@@
Across the three datasets, the Magic Cards consistently exhibited high detection rates, exceeding 97.7% in all configurations, indicating that it is less sensitive to parameter tuning. In contrast, the Lego Sets showed significant variability. For $\text{ef}=200$, the detection rate improved notably with higher values of $M$ (e.g., from 28% at $M=4$ to 30.82% at $M=8$ in $\text{ef}=200$ and $\text{ef}_{\text{construction}}=50$), but beyond $\text{ef}_{\text{construction}}=100$, further increases in $\text{ef}_{\text{construction}}$ provided diminishing returns. This suggests that while denser graphs and more exhaustive index construction improve accuracy for complex datasets, the benefits plateau at a certain point. For the Flowers, the detection rates remained stable at approximately 84.1% across all parameter settings, indicating that this dataset is robust to variations in $M$ and $ef$.
@@ 9 Supplemental Visualization for Conditional Diffusion Models
Figure 6: Examples of the top-25 most influential training data samples for the generated image (the 1-st column) on SD 3 Medium with LoRA, shown from high to low influence from left to right.